OpenStudy (anonymous):

13+ w/7 (fraction)=-18

jimthompson5910 (jim_thompson5910):

The equation is \[\Large 13+\frac{w}{7}=-18\] ???

jimthompson5910 (jim_thompson5910):

or is it \[\Large \frac{13+w}{7}=-18\] ???

OpenStudy (anonymous):

the second one!

jimthompson5910 (jim_thompson5910):

ok thx

jimthompson5910 (jim_thompson5910):

\[\Large \frac{13+w}{7}=-18\] \[\Large 13+w=-18*7\] \[\Large 13+w=-126\] \[\Large w=-126-13\] \[\Large w=-139\]

OpenStudy (anonymous):

-126 = 13 + w w = -126 - 13 = -139